need to see solution to this problem (x+3)*(x+1) = 15; what is X (X=2)
Answer by richwmiller(17219)

worked out it becomes
x^2+4 x+3 = 15
x^2+4 x-12 = 0
(x+6)*(x-2)=0
x=-6 x=2